OpenStudy (anonymous):

If g'(5)=10 and h'(5)= -4, find f''(5) for f(x)=3g(x)-2h(x)+3?

OpenStudy (anonymous):

\[\frac{ df }{ dx }=3*\frac{ d }{ dx }\left[ g(x) \right]-2*\frac{ d }{ dx }\left[ h \left( x \right) \right]+0\] by the sum/difference rules, constant multiple rule, and constant rule for differentiation. This simplifies to \[f \prime(x)=3*g \prime(x)-2*h \prime(x)\] In the same way, \[f \prime \prime(x)=3*g \prime \prime(x)-2*h \prime \prime(x)\] and \[f \prime \prime(5)=3*g \prime \prime(5)-2*h \prime \prime(5)\] Since no information is given on \[g \prime \prime(5)\] and \[h \prime \prime(5)\] I believe this problem cannot be solved. However, if there is a typo, and the problem is actually asking for \[f \prime(5)\] then \[f \prime(5)=3*g \prime(5)-2*h \prime(5)\] \[f \prime(5)=3*10-2*-4\] \[f \prime(5)=38\]